Heat supplied to raise steam in the boiler is 3000 KJ/kg while heat rejected by the system to
the cooling water in the condenser is 2000 KJ/kg. The turbine develops power of 1500 KW and
the work required to pump the condensate back into the boiler is 50 KW. Find the steam mass
flow rate.
1
Expert's answer
2020-06-08T10:28:05-0400

Solution.

"Q_b=3000KJ\/kg;"

"Q_c=2000KJ\/kg;"

"W_t=1500KW;"

"W_p=50KW;"

"Q=Q_b-Q_c;"

"Q=3000KJ\/kg-2000KJ\/kg=1000KJ\/kg;"

"W=W_t-W_p;"

"W=1500KW-50KW=1450KW;"

"\dfrac{W}{Q}=\dfrac{1450KW}{1000KJ/kg}=1.45kg/s;"

Answer:"\dfrac{W}{Q}=1.45kg/s."